OpenStudy (anonymous):

A rectangle has a length 25% longrr than the width, and an area of 80cm squared. Determine the length and width.

OpenStudy (anonymous):

8 and 10

OpenStudy (anonymous):

But how do you solve it

OpenStudy (anonymous):

well 25% longer means l=(5/4)*w and since l*w=80 => (5/4)*w^2 = 80 so w^2=64

OpenStudy (anonymous):

Thannks so much!